We discussed the current academic reserves and circulation enhancement suggestions on Sirsi’s web forum and decided to vote YES for the following, with enthusiasm level and comments indicated in parentheses:
Judy will cast votes by the 11/15/05 deadline. In future the group coordinators will bring new enhancements to ASC for discussion and cast votes for RUL.
A few functional groups still need to report in with revised purpose statements and preferred names. “Team” and “Group” are the top choices so far.
We continue to see users with duplicate barcodes, some not on the list identified by RUconnection. We discussed RUconnection’s suggestion that all students receive new barcodes when cards with student ID numbers are issued beginning in February; we decided it would be too disruptive to our users and not good library public service. We will ask RUconnection to re-issue existing barcodes.
Reminder about using correct mailing addresses for Hungarian and PALCI items. Guidelines for filling RDS and ILL microform requests have been revised. The Securities & Facilities FG will discuss pre-security gate bag checking at Alex and Douglass.
Circulation: The group has been working on revising its purpose statement; researching format and holders for nametags; studying NBL and RUL staff web pages; drafting a revision of RA instructions and application form.
Collection Management: Work continues on transfers to Annex: Chem Abstracts from Chemistry, Lecture Notes in Computer Science from Math, bound periodicals from Physics. Shift of Douglass periodicals will begin the end of November. Andy is conducting sample inventories using Pocket Circ at Chemistry, Art, and Kilmer.
Billing: The Billing group has had preliminary discussions about billing for video (Bookings), including timeframes and notices.
Reserves: Group concluded new file size guidelines work well, monitoring web forums now, ongoing discussions on syllabus formats, evaluating permanent reserve collections at all units, will evaluate RU annual report collection at Alex ugrad reserves.
RDS: Group members have joined Sirsi circulation and demand management listservs and are monitoring web forums. A 137 page paid article request resulted in large file size but presumably it was delivered via the web OK. Powerpoint from RDS training will be revised with a public/user orientation.
Security and Facilities: Two student security workshops were held last month, 1 session was videotaped for sharing. The group may recommend this session become a required part of new student orientation and will coordinate with the Student Coordinators FG’s training. Two members attended a computer security workshop given by OIRT. Bag inspections will be on the group’s next meeting agenda.
Student Coordinators: Reviewing purpose and name. Continuing to monitor use of WebCt, Blackboard, and Sakai on campus, and implications for converting content. Discussing how to prevent unsatisfactory student workers from being reassigned or rehired at another library.
Roger: SERC and Annex positions should be finalized shortly. Two NBL staff will attend METRO workshop on student supervision. Planning for customer service training is underway. Positive comments received on Access Services’ table display and staff participation in State of the Libraries.
Penny: Shifting is on track at Douglass; the Info Services Team is discussing re-naming the rooms (east, west, Tapestry, etc.) when completed. Access staff and some librarians will start wearing nametags after choosing an agreed upon prototype tag; a new calendar system will be in use by the spring semester at Douglass.
Michele: LSM is at full staff; Kevin Conover is the new evening reserve assistant. Weeding is ongoing and the Special Collections Room is filling up with withdrawn books.
Dean: Adam has joined the student and circulation functional groups. Chang exit statistics are up due to new temporary classroom building next to the library. Reports are being used for shelfreading.
Barry: Kilmer is at full staff; Shawn Moriarity is new reserve coordinator and working in Media Services evenings, has Tues-Sat schedule. Daphne Roberts is working some hours at Kilmer’s circulation desk.
Darryl: delivery of new scanner at LSM has been postponed, waiting for Ariel 4.1.1. The link to Imaging Services web page will move up on RUL website, easier to find.
Katie: Shared feedback on videoconferencing functional group meetings— there’s a need for 2 microphones, especially in big groups. Robeson has 3 rooms available for booking meetings: #1 Rm 290, the conference room is the main room, #2 Rm 204 is the room for one person, and the Electronic classroom can also be used but equipment needs to be moved into it. Inventory of Ls and Ms at Robeson is done; the main floor will be complete when Ns are finished.
Glenn: News article on RUL web site reports ILL was a net lender in October, including E-ZBorrow transactions. Personal contact with staff at the British Library is helping to facilitate requests. ILL staff will be attending a Rapid demo this afternoon.
Jeff: Busy at Alexander, with patron issues, damaged items, duplicate barcodes.
Chris: Free Report running time is now available Tues-Thurs., 2-6 p.m. GL3.0 was loaded into the Test system on Monday and while the upgrade went fine, memory parameters need tweaking which require Dave Hoover’s involvement and server configurations. There’s no date yet as to when GL3.0 will be available for testing. If you aren’t receiving regular Unicorn reports, please notify Systems. Chris is studying inventory issues identified by samplers (Ann W. and Andy) and fine-tuning PocketCirc. LIS members attended Sirsi’s remote training for Director’s Station.
NEXT Meeting: Thursday, December 8, 2005, 9:30 a.m. Alexander